OpenStudy (anonymous):

The cost, C, to produce b baseball bats per day is modeled by the function C(b) = 0.06b^2 – 7.2b + 390. What number of bats should be produced to keep costs at a minimum? 27 bats 60 bats 174 bats 390 bats

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \[-\frac{-7.2}{2\times .06}\] whatever that is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

use a calculator

OpenStudy (anonymous):

or don't \[\frac{7.2}{.12}=\frac{720}{12}=60\]

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

yw